# Patterns of Patterns III

### v1

@@ 4. Design Patterns for Pattern-Competent Agents
The PLACARD methodology, as developed and refined across multiple participatory workshops, demonstrates the power of design patterns not only as tools for structuring action, but also as scaffolds for shared sense-making and reflection. Building on this foundation, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns as part of collaborative inquiry. These agents would be expected to:
@@
  • Use flexiformal representations of pattern logic,
  • Operate as a collective, i.e., to form multi-agent systems (MAS)
@@
4.1. Recognizing Pattern-Relevant Situations
The capacity to recognize situations where a pattern might apply lies at the heart of pattern competence. This is not a trivial matching task. It requires the ability to perceive and interpret contextual cues, to identify when a recurring structure or problem-type is emerging, and to discern salience: which aspects of a situation matter, and why. In participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction. When developing artificial agents, the challenge is to develop mechanisms—symbolic, statistical, or hybrid—that allow for similarly situated recognition. Here’s an example showing how this might work for today’s computer systems:
Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously tagged “frustrated” interactions.
Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., a more empathetic tone or an escalation offer).
Next-generation systems might deal with the same scenario in more flexible ways, for instance, not only on the basis of stimulus-response pairs, but more fluidly—for instance, understanding how it was that the customer became frustrated and initiating changes to the company policy accordingly. This more meaningful form of ‘pattern recognition’ requires readiness to see and make sense of structure, and to reframe experience in meaningful, actionable terms — rather than to set data aside as a block, or (which is much the same) to cope with it in a purely reactive way.
4.2. Flexiformal Representation of Pattern Logic
To support agent-based reasoning, patterns must be expressed in a format that is both structured and open-ended. Drawing on the modified pattern structure used in PLACARD—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
@@ 4.2. Flexiformal Representation of Pattern Logic
a broad frame (task domain, discourse setting) that activates relevant subsystems,
@@
a salient event or state within that frame,
@@
a tension or constraint that introduces conflict,
@@
While logic and reasoning have been core topics in computer science from its outset (Sowa, 1999), further research is needed to determine what sorts of reasoning—clearly combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans to think about patterns.
For now, we refer to the overall approach as “flexiformal” (Kohlhase, 2012), because this leaves the door open to combining different formal systems in a suitably flexible fashion. Simply adding more patterns won’t be enough to implement human-like reasoning in pattern-capable systems. For one thing, the patterns need to be used to dynamically allocate cognitive resources.
4.3. Reflection Loops and the Evolution of Epistemic Infrastructure
To sustain meaningful pattern use, agents must also be able to reflect. This means embedding meta-level structures like the PAR into the environment itself. Reflection loops should support the capture of learning, the revision of pattern catalogues, and the emergence of new design insights.
In future architectures, agents and humans alike can annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence. We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).
@@ 4.4. Rethinking the Pattern Catalogue
The PLACARD A/B/C pattern catalogue was created with human participants in mind. It emphasizes social affordances, embodied prompts, and facilitation moves. While some of these translate readily into multi-agent contexts—e.g., “Meta-Reflection Loop” or “Pattern the Play”—others rely on human sensory input or informal framing strategies that are not directly available to AI systems.
Thus, we propose a tripartite model of pattern types needed for pattern-competent agents, generalizing the layers we used previously.
1. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
PAR structures collaborative reflection and coordination. Interactional patterns in general guide agents in playing roles, reflecting, and negotiating meaning in collaborative settings—essentially the micro-level behaviors that mirror PAR’s inquiry loop. Examples include:
  • Role Reveal: Adopt explicit roles in multi-agent dialogue.
@@
These patterns require some translation to make it understandable to the machine, e.g., encoding roles as functions in MAS architectures or structuring turn-taking rules.
2. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
CLA unpacks multiple layers of meaning—litany, system, worldview, myth. These cognitive patterns help agents reason across contexts and detect tensions, embedding depth and nuance in internal reasoning. They are interpretive structures, supporting internal reasoning and pattern use:
@@
3. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
DPL captures reusable structures for action. These infrastructural patterns define persistent templates and system-wide protocols that support reuse, feedback, and evolution of the broader system design. For example these are 3 meta-patterns that could enable engagement with patterns:
  • Shared Memory Architecture: Persistent storage of pattern use and modification. Putting patterns where people can get them, like the catalogue, akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015)
@@
They govern the architecture of the agent system much like DPL governs the architecture of human practice.
4.5. Desiging Environments for pattern-competent MAS
In addition to broader contextual information, relevant multi-agent system (MAS) environments would include the social and cognitive dynamics of inquiry: coordination, negotiation, reflection. Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005). This distributed adaptation mirrors cultural evolution: strategies become shared, revised, and recontextualized. Rather than simply evaluating success in terms of task completion or efficiency, MAS environments invite questions about “how” solutions are created.
This reflection constitutes suggests a key amendment to PLACARD. Whereas PLACARD was described in classic psychological terms (Hilgard, 1980) — which are also taken over by roboticists (Cassimatis et al., 2004) — it does not specifically talk about the environment in which the agents it considers would be situated. In robot terms, for example, we would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning. This way of thinking suggests that the PLACARD pattern language should be extended to be an “A/B/C/D catalogue”, where the fourth section contains environmental or developmental patterns. With this in mind, in addition to further detail on the candidate patterns for pattern-competent agents — as outlined above — the appendix also includes some patterns that would guide the design of their working environments.
5. Virtual Multi-Agent Workshops in the Spirit of PLACARD
To explore the idea of pattern-competent agents developed Section 4, we conducted a series of virtual workshops using fictional agents within ChatGPT.  Using the PLACARD method, those agents rotated through breakout sessions, engaged in pattern-guided inquiry, and made use of a curated and emergent pattern catalogue. It is worth emphasizing that the workshops were computer-aided design fictions (we have previously explored this way of working “by hand” (Danoff et al., 2019), and we’re aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025)). ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns, but it cannot build and run custom agent architectures. All ‘agents’ were textual personas that were simulated by one model. The agents were plausibly able to discuss the design patterns and employ them in light-weight design exercises. Below is an excerpt from one such workshop, in which several virtual agents begin collaboratively applying the PLACARD methodology to plan a research project on epistemic AI. The agents have somewhat stereotyped fictional biographies which to some extent shape their responses (e.g., “Carlos – Epistemology & Systems Philosophy”; “Mei – Participatory Design & Feminist STS”).
Excerpt: PLACARD Planning Dialogue on Epistemic AI Carlos: I propose we begin with CONTEXT SETTING. If we’re designing a research project on epistemic AI, we need a shared sense of what “epistemic” means in our context. Priya: Agreed. For me, it’s about how agents—human or artificial—come to know, not just calculate. So I’d emphasize design patterns that support interpretive, situated reasoning. Raj: Right. Maybe we could use Borrow the Situation—look at real-world cases of failure or success in knowledge systems. Misinformation management? Scientific discovery pipelines? Sofia: I think we need to Name the Frame explicitly. Are we researching tools for knowledge validation? Or institutional models for human-AI epistemic partnerships? Mei: Maybe both. What if we used Pattern the Play to design speculative workshop sessions? Each session could simulate a different epistemic tension—uncertainty, disagreement, novelty—and trial agent responses. Yuki: And we need some kind of Meta-Reflection Loop. A structure where, after each workshop, we reflect not just on what we learned, but how the patterns performed. This could feed into our developmental arc. Carlos: So the goal of the series isn’t just content generation—it’s to test, refine, and expand the pattern catalogue itself? Sofia: Exactly. And that aligns with the DO YOUR RESEARCH pattern: we’re generating grounded material and modeling the tools that structure knowledge formation. Raj: We’ll need a pattern-aware toolchain too—maybe a Pattern Knowledge Repository that evolves with each workshop. Otherwise we lose coherence. Priya: Let’s add a Timebox the Core phase to every session. This gives us just enough structure to avoid sprawling, without over-engineering. Mei: I’ll draft a plan using PLACARD as a scaffolding template. We can each take ownership of one layer—PAR, CLA, DPL—and develop sub-patterns and protocols that fit.

### v2

@@ 5. Design Patterns for Pattern-Competent Agents
The typical use of design patterns is as methods for structuring action. As the previous section illustrates, they can also be used to scaffold sense-making and reflection after the fact. Building on these foundations, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns in a process of collaborative inquiry. These agents would be expected to:
@@
  • Use suitably flexible representations of pattern logic,
@@
  • Operate as a collective, i.e., agents combine in multi-agent systems (MAS),
Our existing patterns give a start in this direction, at least insofar as they could help us to imagine collaborative workshops that include or are populated by artificial agents. However, the aim of embedding the abilities mentioned above within the computational agents will suggest some new patterns. First, let’s consider the above points in turn.
5.1. Overview of requirements
5.1.1. Recognizing and Responding to Pattern-Relevant Situations
The capacity to recognize situations where a pattern might apply lies at the heart of pattern competence. This is not a trivial matching task. It requires the ability to perceive and interpret contextual cues, to identify when a recurring structure or problem-type is emerging, and to discern salience: which aspects of a situation matter, and why. In typical participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction. For people who are used to working with design patterns, a new experience might be thought with some light attention going to an existing pattern catalogue, without being constrained to only think in terms of known patterns. When developing artificial agents, the challenge is to develop mechanisms—symbolic, statistical, or hybrid—that allow for similarly situated recognition.
Here’s an example showing how this sort of process works for today’s computer systems:
Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously interactions that were tagged “frustrated”.
Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., to adopt a more empathetic tone or make an escalation offer).
Next-generation systems might deal with the same scenario in more flexible ways, for instance, by attempting to build a narrative understanding of how it was that the customer became frustrated, and proposing changes to the company policy that would avoid such problems in the future.
This more meaningful form of ‘pattern recognition’—which is set up in a way that supports ‘pattern application’—requires readiness to see and make sense of structure, and to reframe experience in appropriately meaningful, actionable terms.
5.1.2. Flexiformal Representation of Pattern Logic
The primary difference between design patterns and computer programs is that design patterns are both structured and significantly open-ended. Their structure comes primarily from the use of a consistent schema (e.g., context, problem, solution), and from links between patterns. Their open-endedness lies in their reliance on user interpretation and situational judgment. Insofar as they ‘operate’ at all, patterns do so through enactment in specific contexts, rather than via deterministic execution. Nevertheless, as we explore further here—and as been discussed by various authors before (Corneli et al. (2018) and references therein, particularly to the work of T. P. Moran and Rivka Oxman)—patterns can also be framed in computational terms.
Drawing on the adapted pattern template we used to describe our PLACARD patterns—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
@@ 5.1.2. Flexiformal Representation of Pattern Logic
A broad frame—task domain, discourse setting, data ingest—that activates relevant subsystems.
@@
A salient event or state within that frame, which can be selected by preliminary processing, such as filtering or querying.
@@
A tension, constraint or conflict—established through additional filtering, querying, data alignment—which suggests the possibility of, or need for, further processing.
@@
A set of critical checks or tests to take before selecting the action; typically framed with respect to some existing body of empirical evidence; possibly also useful for supplying auditable reasoning alongside the action.
For now, we refer to the overall approach as “flexiformal” (Kohlhase, 2012), because this leaves the door open to combining different formal systems in a suitably flexible fashion. Logic and reasoning have been core topics in computer science from its outset (Sowa, 1999); however further research would needed to determine what sorts of reasoning—pesumably, combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans when they think about patterns. Furthermore, simply specifying more patterns won’t be enough to implement human-like reasoning in pattern-capable systems. As we’ve indicated here, a key thing that the patterns need to do is to dynamically allocating cognitive resources. The patterns’ internal structure and external relationships with each other (and with other interfaces) constrain the system’s ability in that regard.
5.1.3. Reflection Loops and the Evolution of Epistemic Infrastructure
To sustain meaningful pattern use, agents would have to be able to reflect; noting new patterns, revising them, including changing the way they relate, and deprecating or deemphasising patterns that are no longer relevant. This requires embedding meta-level review structures—perhaps something like the PAR—within the agents or their interactions. In future architectures, agents and humans alike might work together to annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence. (We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).)
5.1.4. Environments in which pattern-capable agents can form pattern-competent MAS
Relevant multi-agent system (MAS) environmental features would include scaffolding for the social and cognitive dynamics of inquiry, such as coordination, negotiation, and reflection (as touched on above). Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005). Rather than simply evaluating success in terms of task completion or efficiency, pattern-competent MAS environments might invite questions about “how” solutions are created.
@@ 5.2. Rethinking the Pattern Catalogue
The series of short reflections above suggests an extension to the way we were thinking about the PLACARD pattern. PLACARD was cast along classic psychological lines (Hilgard, 1980)—and the sensory/cognitive/motor facilities have also been taken over by roboticists (Cassimatis et al., 2004). However, it does not specifically talk about the environment in which the agents it considers would be situated. We would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning; and, perhaps even more to the point, we would expect a robot that operates as part of a swarm to work very differently from one that operates individually. Perhaps because the default assumption with PLACARD was that the actors were humans operating in standard human ways, it did not deal with these contextual issues in any detail.
Accordingly, the four dimensions outlined above suggests that a pattern language for pattern capable agents should be extended—retaining the A. sensory, B. cognitive, and C. motor segments, and adding a fourth segment, Denvironmental or developmental patterns.
With this reflection in mind, we propose a quadripartite model of pattern types needed for pattern-competent agents, generalizing and extending the layers we used previously, and giving further expression to the requirements outlined above. Some examples follow (detailed pattern statements are in the Appendix).
A. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
Interactional patterns in general guide agents in noticing when patterns can be applied, playing roles, reflecting, and negotiating meaning in collaborative settings. To a reasonable extent these behaviors mirror PAR’s inquiry loop, but instead of being a stand-alone reflection ritual, they would describe a broader set of practices of mutuality within an environment. Examples could include.
@@
  • Role Reveal: Adopt and express explicit roles to guide multi-agent dialogue.
B. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
These cognitive patterns help agents reason across contexts and detect tensions, embedding depth and nuance in internal reasoning. Whereas CLA unpacks multiple layers of meaning, these patterns describe further interpretive structures, supporting internal reasoning and pattern use:
@@
C. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
These infrastructural patterns define templates and system-wide protocols that support various forms of interaction with and through the broader system—reuse, feedback, and evolution of the broader system design. Whereas DPL captures general structures for action, these patterns focus on the self-coherence implied by ‘action’, insofar as it always involves an actor. For example, we begin here with three meta-patterns that would enable agents to engage with a collection patterns:
  • Shared Memory Architecture: Persistent storage of pattern use and modification, putting patterns that aren’t being actively used someplace where they can be retrieved (akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015))
@@
D. Environmental or Developmental patterns (e.g., institutional arrangements)
These patterns help multi-agent systems evolve over time and maintain coherence across shifting contexts. They provide orientation at the level of institutions, ecological awareness, and collective change.
  • Institutional Drift: When pattern use becomes routine but out of sync with current challenges, trigger a meta-level review to restore relevance.
  • Norm Seed: At the start of a new MAS, plant simple, precedent-based mini-patterns to guide initial interactions and shape future dynamics.
  • Boundary Oscillation: When divergent roles or norms emerge across domains, activate a boundary negotiation process to realign shared pattern use.
  • Reflection Scaffold Upgrade: If reflection fails to generate insight, introduce new layered tools or protocols to support deeper system learning.
6. Virtual Multi-Agent Workshops with Patterns
To explore the idea of pattern-competent agents developed in Section 5, we conducted a series of virtual workshops using fictional agents within ChatGPT; we report on two such workshops here.
6.1. Simulated PLACARD workshop
Agents were invited to use the PAR, CLA, and DPL methods, and were given access to existing patterns from our previous work with PLACARD, along with a further collection of new patterns covering similar topics (Appendix:(1)). The prompt for the workshop appears in Figure 2, edited for readability. This adapts the prompt template from Section 3.3. The agents that were created were given somewhat stereotyped fictional biographies, which to seem to shape their contributions to the dialogue. However, ChatGPT cannot build and run custom agent architectures. All ‘agents’ were textual personae that were simulated by one model.
Virtual workshops in this style should be understood from the outset as computer-aided design fictions. ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns. In this example, it would be most accurate to say that a virtual workshop was “simulated”, rather that that it “actually happened”.
Some key experts from the ChatGPT’s trace (obtained via the Think Deeply feature) are as follows:
I’m gathering key workshop patterns and considering roles like “Functional Roles” and “Facilitator Roles” to ensure synergy and leverage interdisciplinary insights.
Wondering about talents like Omar, focusing on knowledge representation, and Yuki, adding a creative vision to multi-agent interaction. Considering resources and striking a balance.
OK, let’s see. Participants are thinking through research questions in breakout rooms, covering math, multi-agent, and meta topics, ensuring a balanced mix of expertise.
Participants are circling back old and new patterns, making iterations, and planning the final steps. Pattern names stand out, guiding their approach for a comprehensive strategy.
See Figure 3 for a summary of the workflow, and main research questions that were considered. This figure summarises a considerably longer (22-page) workshop report that was generated at the end of the run. [^1] https://bit.ly/virtual-placard-workshop. The linked supplemental information includes the verbatim prompt which is glossed in Figure 1.
Relative to our discussion of patterns: looking over the report, existing patterns (like REINFUSE EXPERTISE, GOING META, Start with Sense, Name the Frame, etc.) have been applied in a plausable way. Some new proto-patterns e.g., (“Lab Shadowing” and “Checklist Manifesto for AI”) were also proposed.
Note that the high-level themes in the workflow—the breakout rooms—are derived from earlier chats with ChatGPT. While the methods we have described are repeatable, but it would probably be impossible to reproduce the experiment exactly, even using the same ChatGPT account (particularly because other chats have taken place using that account in the mean time).
ChatGPT is able to generate plausable dialogue snippets on request. If it had asked for “a detailed simulation of the breakout session” as an output, or immediately after running the workshop simulation, it would likely produce something plausible. In the following section, we did just that.
6.2. Simulated pattern-capable agents workshop
Reflecting on the strengths and shortcomings of the virtual workshop described above, we tried another design. The prompt is given in Figure 4. This time, the collection of patterns for pattern-capable agents and environments (in Appendix(2), Appendix(3)) played a key role in the prompt.
6.2.1. First round
With Figure 4 as the overall framing, the subprompt and short answer for this round was:
Q1. “How would we characterize, in prose, the constraints and features that the agent-environment patterns supply for collective problem solving?”
A1. “Each pattern-agent imposes certain constraints that structure the collective problem-solving process (preventing particular failures or excesses), and each provides features that enable or enhance the group’s ability to solve problems together.”
Full outputs are available as a supplement. [^2] https://bit.ly/virtual-agents-workshop-i The following changes to the pattern catalogue were proposed on the way to the short answer above.
EVOLVE – Timebox the Core: The Reflect in Layers and Timebox the Core agents collaboratively proposed a change to the Timebox the Core pattern. The pattern was evolved to include a provision for deeper reflection: when core tasks are finished or when a deeper insight is flagged during the session, Timebox will allow scheduling of a follow-up for that deeper exploration rather than strictly ending the session. This evolution was agreed upon by the group and validated via Legitimate Iteration for consensus and proper documentation.
GENERATE – Outcome Synthesizer: The agents identified a gap – no existing pattern solely focused on integrating multiple contributions into a coherent solution. Meta-Reflection Loop (supported by Candidate Move Generation and others) proposed a new pattern-agent, tentatively named Outcome Synthesizer. With encouragement from Self-Patterning Mandate and formal group approval (again facilitated by Legitimate Iteration), this new pattern was introduced. Outcome Synthesizer is designed to gather and synthesize the group’s disparate inputs into a unified outcome or summary, ensuring the collective output is coherent.
(Note that the Outcome Synthesizer is effectively a rediscovery of the WRAPPER peeragogy pattern (Corneli et al., 2015)).)
6.2.2. Second round
This time, we used the following subprompt, and received the accompanying short answer: [^3] https://bit.ly/virtual-agents-workshop-ii
Q2. “How would the pattern-capable agents take over the work from the previous PLACARD workshop on epistemic AI, as captured in a printout describing the same, and build on it? Please phrase the answer as a prompt for another round of agent-based simulation.”
A2. “How might we run a workshop where AI Fellows co-develop new institutional forms for epistemic AI research, supported by our pattern-agent infrastructure?”
In addition to “breaking the fourth wall” and proposing to run a workshop with other humans (or simulated humans?), something else interesting and a bit strange happened in the dialogue itself. The agent workshop started to treat some of the PLACARD patterns that the system had encountered previously as ‘agents’, effectively drawing them into the dialogue. One outcome of the workshop was to propose an evolution to Increase Participant Control—even though this was not one of the patterns that had been submitted for revision.
While the answer A2 could be interesting to adapt and explore in a further workshop—and it is indeed quite intriguing to think about setting up a hybrid workshop with some agents and some humans via a chat interface [^4] https://github.com/IngoKl/PolyBotConversation—it is perhaps best to close the curtain on this series of AI workshops, and discuss what we have seen so far.
“Remember how we did some simple multi-agent workshops before? Let’s simulate the architecture described in our paper-in-progress with some ChatGPT-based agents. We are going to base the agents on a diverse group of Research Fellows in AI. Let’s consider 10 such agents. Please additionally include a Recorder who will keep track of what happens in which workshop session. Please also include a Reporter who can speak to what happens. The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, which are available in the file patterns.txt. The outcome of the series should be a detailed plan of work for a research project on epistemic AI. You should do deep research to make sure that the discussions are plausible.”
Figure 2. Prompt Given to Generative AI Program to create and run a Virtual Workshop (lightly edited for readability)
Round 1: Exploring Key Research Questions (Ideation Round)
  Breakout room 1 (Epistemic Modeling & Validation in Mathematical Domains): (Q1) What frameworks can rigorously verify AI-generated proofs and integrate with human mathematicians’ validation workflows? and (Q2) How can we model the epistemic confidence in a mathematical claim across human–AI collaborations?
  Breakout room 2 (Collaborative Design Patterns for Multi-Agent Systems): (Q3) What design patterns can facilitate transparent communication and division of labor in mixed teams of humans and AI agents? and (Q4) Which collaborative strategies best mitigate conflicts or errors in autonomous multi-agent research systems?
  Breakout room 3 (Metascientific Observation of AI Research Practice): (Q5) What methods can we use to observe and measure current AI research practices (in labs and publications) to identify gaps in transparency and collaboration? and (Q6) Which interventions (patterns, policies, or tools) could improve reproducibility and inclusivity in AI research communities?
Coffee break
  “Early cross-pollination: Fellows are already trading metaphors across domains – a sign of interdisciplinary synergy.” PAR
  “I realized my ‘big question’ about AI proofs was actually two questions once I considered both technical and human angles.”
Round 2: Developing Approaches and Patterns (Design Round)
PAR “We’re actually developing our own pattern for how to run a project like this – maybe we should write that down too.” SHARE BACK plenary
Round 3: Integration and Planning (Convergence Round)
Integration and Closing Discussion
Figure 3. Reported workshop structure and selected quotes
You have been given 20 design patterns describing pattern-capable agents and pattern-capable agent environments. I want you to run a multi-agent textual simulation where each pattern is an agent. “Agents” will operate in an environment which has the agent-environment patterns as constraints or features.
You are tasked with the following problem:
We need to design a multi-agent protocol that can address real-world problems. There are 20 of us now in total. We can create more agents as needed. Each ‘agent’ is a pattern, and each pattern is an ‘agent’ — including the agent-environment patterns. Together, we are to generate sub-problems that elaborate this challenge and work through them in a self-organised way.
In addition to “DISCUSS”, our two main moves are:
  • EVOLVE: Revise or adapt the description of an existing pattern (including changing its trigger conditions, actions, or connections).
  • GENERATE: Create a new pattern-agent who will join the population, based on observed needs or tensions.
You may also EVOLVE or GENERATE agent-environment patterns. These define the operational constraints or affordances within which all agents must function. Changes to these patterns will be treated as systemic shifts, impacting the context for subsequent rounds.
We will, interactively, run several rounds of simulation. Each round will consist of:
  1. (1)
    Agreeing upon a sub-prompt or tension for the group to address.
  2. (2)
    Discussion between the agents.
  3. (3)
    One or more agents making EVOLVE or GENERATE moves in response.
  4. (4)
    Optional meta-commentary or reflection from other agents.
At the end, output a heredity tree showing how the agent population evolved — i.e., which agents were changed, what was added, and what led to each transformation.
Please include:
  • A summary of the dialogue for each round
  • A clear record of EVOLVE / GENERATE moves
  • The updated pattern descriptions (for any evolved or new agents)
  • A final heredity tree visualising the evolution
Figure 4. Prompt for evolving the patterns for pattern-capable agents and agent-environments